OpenStudy (anonymous):

write a polynomial in standard form with the following zeros: -3, 2, and 5

OpenStudy (imstuck):

First let's put them into their factorizations. Those numbers alone mean that x = -3, x = 2 and x = 5. But if we work backwards from factoring we would have (x + 3) = 0, (x - 2) = 0, and (x - 5) = 0. Those are the roots now of the polynomial. Working backwards still we have (x + 3)(x - 2)(x - 5). Multiply all those together by FOIL-ing a couple of times and you will have you answer. BTW, you might should know that if a polynomial has 3 roots, it is an x^3 equation. The number of roots a polynomial has defines the degree of the polynomial. Any other questions on this or anything else, feel free to ask!

OpenStudy (anonymous):

thank you so much! im going to try to solve it and then ill tell you my answer. when i send it, could you tell me if its right?

hero (hero):

If -3, 2, and 5 are the zeroes of the polynomial, then x = -3 x = 2 x = 5 Moving everything to the left side: x + 3 = 0 x - 2 = 0 x - 5 = 0 By zero Product Property (x + 3)(x - 2)(x - 5) = 0 After multiplying you end up with a cubic polynomial f(x) = (x + 3)(x - 2)(x - 5)

OpenStudy (anonymous):

hero, is f(x) = (x+3)(x-2)(x-5) the polynomial in standard form? or do i need to foil this to get to the standard form? if so, how do you foil with three terms?

hero (hero):

Use the distributive property rather than FOIL. This is an anti-FOIL community: f(x) = (x(x - 2) + 3(x - 2))(x - 5)

OpenStudy (anonymous):

where did you get the 3 on the outside of (x-2) and where did you get the x on the outside of (x-2) ?

hero (hero):

From x + 3... I'm sure you remember the distributive property: (b + c)a = ba + ca In this case, when you distribute (x + 3)(x - 2), you treat (x - 2) as \(a\) , then \(x\) as \(b\) and \(3\) as \(c\) So (x + 3)(x - 2) = x(x - 2) + 3(x - 2)

OpenStudy (anonymous):

so, would the answer be, x squared +x-6

hero (hero):

(x + 3)(x - 2) = x(x - 2) + 3(x - 2) = x^2 - 2x + 3x - 6 = x^2 + x - 6

hero (hero):

But now you have to perform the same distributive process for (x^2 + x - 6)(x - 5) But this time, when you do it, show your work.

OpenStudy (anonymous):

ok thank you ill send you my answer when im done

hero (hero):

Along with your steps. I'm more interested in your steps than the answer.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

x(x^3+x-6) -5(x^2+x-6) distribute the x to 1st parenthesis, -5 to the second parenthesis x^3+x^2-6x-5x^2-5x+30 combine like terms x^3-4x^2-11x+30 the final answer

hero (hero):

Interesting, you made a slight mistake at the beginning. You wrote: x(x^3 + x - 6) - 5(x^2 + x - 6) But you meant to write: x(x^2 + x - 6) - 5(x^2 + x - 6)

OpenStudy (anonymous):

oh ok i see where i did that. on my paper its copied down right. so thanks :) can i ask another math question?

OpenStudy (anonymous):

if yes, here it is: Divide and express the result in the form P(x) = D(x)Q(X) + R(X) (4x^3 + 2x^2 + 3x + 4) divided by (x+4)

hero (hero):

@beachbum1996, you still here?